Which fraction is equal to:







a. a quarter







b. a sixth and







c. two thirds , respectively


1
Expert's answer
2022-02-23T09:24:43-0500

a) The fraction of a quarter is equal to "\frac{1}4."


b) The fraction of a sixth is equal to

"\frac{1}6."


c) The fraction of two thirds is equal to "\frac{2}3."